# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

### v3

@@ 2.2 Control Barrier Functions
Lemma 2 implies that a safe control input on the boundary of the set $S$ can render the trajectory $\mathrm{x}$ forward invariant in $S$. Control barrier functions (CBFs), first introduced in [4], additionally impose conditions on the input when the trajectory is strictly inside the set before reaching the boundary, which enables the trajectories to “smoothly brake” as they approach the boundary. Here, we extend its definition to robust CBF for the dynamics with disturbance (8).
@@
A function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Assumption 3 for a closed set $S$ is a robust CBF for the dynamics (8) if there exists an extended class $\mathcal{K}$ function $\alpha$ such that, for all $x\in S$,
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)+\alpha\left(h_{S}(x)\right)\geq 0.$ (11)
@@
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)+\gamma h_{S}(x)\geq 0.$ (12)
This is the most common choice of class $\mathcal{K}$ function used in the CBF literature, and enables us to make a connection between CBFs and reachability value functions where $\gamma$ will play the role of a discount factor in the reachability formulation. Intuitively, (12) ensures that $h_{S}(\mathrm{x}(t))$ does not decay faster than the exponentially decaying curve $\dot{h}_{S}(\mathrm{x}(t))=-\gamma h_{S}(\mathrm{x}(t))$. This induces the braking mechanism to any trajectory $\mathrm{x}$ approaching the boundary of $S$. If (12) is satisfied, (10) is trivially satisfied at $x\in\partial S$ where $h_{S}(x)=0$. Thus, according to Lemma 3, the existence of the robust CBF $h_{S}$ is a sufficient condition for $S$ being robust control invariant.
@@
Let the dynamics (8) satisfy Assumption 4 and let $S\subset\mathbb{R}^{n}$ satisfy Assumption 2. If a robust CBF $h_{S}$ exists, $S$ is robustly control invariant.
@@
For systems without disturbance (1), if (6) in Lemma 2 holds with strict inequality for a compact and control invariant set $S$, any $h_{S}$ satisfying Assumption 3 is a CBF with large enough $\gamma$ [4, Lemma 2].
@@ 3 Forward Reachability Analysis
In this section, we apply forward reachability analysis to control invariant sets. In order to do so, we first provide background on the forward reachability of a set and introduce definitions of forward reachable tubes.
@@ 3.1 Forward Reachability
Forward reachability analyzes a set’s evolution in the future; its purpose is to identify the states that trajectories from an initial set $C\subset\mathbb{R}^{n}$ reach forward in time. The forward reachable tube (FRT) of a set, roughly speaking, encompasses states that are reached by trajectories that depart from the initial set. This concept is illustrated in Figure 1. For autonomous systems (e.g. $\dot{\mathrm{x}}(t)=f(\mathrm{x}(t))$), the forward evolution of a set is uniquely determined (Fig 1 (a)). However, for systems with control and/or disturbance inputs like (1) and (8), the trajectory, and thus the forward reachable tube, can be determined in various ways according to control and disturbance.
Consider the dynamics without disturbance, satisfying (1). At one extreme, the control can use its best effort to get further away from the original set $C$, and at the other extreme, the control works to stay as close to $C$ as possible. The former would cause the FRT to expand maximally covering all the states such that from $C$, reaching them is viable (Fig 1 (b)), and the latter would induce the FRT to grow minimally, encapsulating only the states which inevitably must have evolved from $C$ (Fig 1 (c)). From this intuition, we are able to define the viable (maximal) and inevitable (minimal) FRTs of a set $C$. Here, we only define the inevitable FRT which is the focus of this study, and readers are referred to [27] for the definition of the viable FRT.
Applications of forward reachability in the safe control and verification literature primarily focused on determining the viable FRT of the set of initial states, and checking whether this set intersects with the unsafe set [3, 23, 37]. In this context, the utility of inevitable FRTs would be limited, since the trajectories from the initial set could still enter the unsafe region even if the inevitable FRT does not [27]. In this paper, we revisit the utility of the inevitable FRT under the specific context in which the initial set is robustly control invariant.
@@
where $\mathrm{u}^{-}:(-\infty,0]\rightarrow U$ is an element of $\mathcal{U}^{-}$, a set of measurable backward control signal. With $\mathcal{D}^{-}$ denoting a set of measurable backward disturbance signals $\mathrm{d}^{-}:(-\infty,0]\rightarrow D$, $\xi_{d}^{-}$ is a non-anticipative strategy for the disturbance backward in time. The separate notation is necessary due to the causality of the disturbance strategy being reversed in time. We will call $\mathrm{x}^{-}$ the backward trajectory of (terminal state) $x$. Evaluating whether $\mathrm{x}^{-}$ reaches the set $C$ when time is considered to flow backward will tell us whether $x$ belongs to the forward reachable tube of $C$.
@@
Note that  $\mathrm{FRT}(\cdot)$ can be interpreted as a set mapping, $\text{FRT}:2^{\mathbb{R}^{n}}\rightarrow 2^{\mathbb{R}^{n}}$. In words, $\mathrm{FRT}(C)$ is a collection of states such that every trajectory reaching it forward in time must have passed through $C$ at some point in the past. The FRT is shaped by the control aiming to restrain the growth of the FRT, whereas the disturbance is assumed to act adversarially and attempts to grow the FRT. For control systems, the inevitable FRT is minimal [27] since it excludes any state that can be reached by a trajectory that does not evolve from $C$, which is particularly relevant to the concept of control invariance.
@@ 3.2 Forward Reachable Tubes of Control Invariant Sets
We now investigate the relationship between the inevitable FRTs and control invariance. The main theorem of the section is as follows:
@@
Suppose $f$ satisfies Assumption 4. Then, a set $S$ satisfying Assumption 2 is robustly control invariant under (8) if and only if $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$.
The theorem identifies that the interior of any control invariant set with a differentiable boundary is a fixed point of $\mathrm{FRT}(\cdot)$. The non-triviality of the theorem arises from the need for not only control invariance but also the differentiability of the boundary of the set $S$.
Note that, in contrast to the forward reachability result in Theorem 1, backward reachability characterizes closed robust control invariant sets as fixed points of the viability kernel operation [5], without requiring the invariant set to have a differentiable boundary. Nevertheless, our motivation for focusing on forward reachability lies in its potential to reveal connections between reachability value functions and CBFs—a connection that will be explored in the subsequent sections. We will show that backward reachability value functions are fundamentally distinct from CBFs, limiting their utility for safe control synthesis.
The rest of this section presents the proof of Theorem 1. For the proof, we have to reason about the backward flow of the dynamics, as in (13). Thus, we first consider the notion of backward control invariance, the mirrored version of the forward control invariance.
Definition 6 (Robustly backward control invariant).
Corollary 1.
Under Assumptions 2, 3, and 4, $S$ is robustly backward control invariant if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{s.t.}~{}-\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D.\vspace{-1em}$ (15)
Figure 2: Forward reachable tubes under double integrator dynamics for various shapes of $S$. In the first row, the initial set $S$ is visualized as the interior of the pink level curve. The interior of the blue level curve in the second row is $\mathrm{FRT}(\mathrm{Int}(S))$ for each case. (a) Smooth $S_{a}$ that is not control invariant, resulting in $\mathrm{FRT}(\mathrm{Int}(S_{a}))\neq\mathrm{Int}(S_{a})$. The FRT is still not control invariant. (b, c) Nonsmooth sets $S_{b}$, $S_{c}$ that are control invariant; in case (b), $\mathrm{FRT}(\mathrm{Int}(S_{b}))\neq\mathrm{Int}(S_{b})$, and in case (c), $\mathrm{FRT}(\mathrm{Int}(S_{c}))=\mathrm{Int}(S_{c})$. This shows that Assumption 2 is required for Theorem 1 to hold. (d) Smooth control invariant $S_{d}$ that results in $\mathrm{FRT}(\mathrm{Int}(S_{d}))=\mathrm{Int}(S_{d})$ according to Theorem 1.
Lemma 4.
Under Assumptions 2, 3, and 4, $S$ is robustly forward control invariant if and only if $\mathrm{Int}(S)^{c}$ is robustly backward control invariant.
Proof.
By Lemma 3, $S$ is robustly forward control invariant if and only if for all $x\in\partial S$, (10) is satisfied. Note that $\partial S=\partial\mathrm{Int}(S)^{c}$ and $\mathrm{Int}(S)^{c}$ and $h_{\mathrm{Int}(S)^{c}}:=-h_{S}$ also satisfies Assumptions 2 and 3, respectively. Since $\frac{\partial h_{\mathrm{Int}(S)^{c}}}{\partial x}(x)=-\frac{\partial h_{S}}{\partial x}(x)$, (10) is equivalent to
In Lemma 4, Assumption 2 guarantees that, for a state $x_{1}$ on the boundary of $S$, if there exists a particular control $u_{1}$ such that $f(x_{1},u_{1},d)$ points inward to $S$ for all $d\in D$, $-f(x_{1},u_{1},d)$ points outwards to $S$ for all $d\in D$.
Definition 7.
[…]
This definition applies the concept of leadership kernel in [12] to the backward dynamics. Given any closed set $C\subset\mathbb{R}^{n}$, the $\mathrm{VK}^{-}(C)$ is the largest negatively robustly invariant set in $C$. Thus, if $C$ is negatively robustly control invariant, $C$ is identical to $\mathrm{VK}^{-}(C)$:
Lemma 5.
A closed set $C\subset\mathbb{R}^{n}$ is negatively robustly control invariant under (8) if and only if $\mathrm{VK}^{-}(C)=C$.
Proof.
This is a direct results of the definitions. ∎
Lemma 6.
$\{\mathrm{FRT}(C)\}^{c}=\mathrm{VK}^{-}(C^{c}).\vspace{-0.5em}$ (17)
Proof.
The complement relationship (17) follows directly from Definitions 5 and 7. That $\mathrm{VK}^{-}(C)$ is a closed set is proven in [12], and $\mathrm{FRT}(C)$ being an open set results from the complement relationship. ∎
Based on the lemmas above, we present the proof of Theorem 1 below.
Proof.
(Theorem 1) $S$ is robustly forward control invariant if and only if $\mathrm{Int}(S)^{c}$ is robustly backward control invariant, by Lemma 4. By Lemma 5, $\mathrm{Int}(S)^{c}$ is robustly backward control invariant if and only if $\mathrm{VK}^{-}(\mathrm{Int}(S)^{c})=\mathrm{Int}(S)^{c}$. By Lemma 6, $\mathrm{VK}^{-}(\mathrm{Int}(S)^{c})=\mathrm{FRT}(\mathrm{Int}(S))^{c}$. Thus, from the above statements, $S$ is robustly forward control invariant if and only if $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$. ∎
The differentiability of the boundary (Assumption 2) is necessary for the crucial step of the theorem’s proof, which uses the equivalence between $f(x,\bar{\pi}(x))\in T_{S}(x)$ and $-f(x,\bar{\pi}(x))\in T_{\mathrm{Int}(S)^{c}}(x)$ (Lemma 4). At the non-smooth boundary of $S$, this is not necessarily true. In this case, a forward trajectory can be inevitably “leaked” from the interior of $S$, leading to the expansion of $\mathrm{FRT}(\mathrm{Int}(S))$ to a strict superset of $\mathrm{Int}(S)$. An example of this incident is introduced next (Figure 2b).
3.3 Example: Double Integrator
We introduce examples of four different initial sets (Figure 2 first row) and their FRTs (Figure 2 second row) to illustrate the importance of two key assumptions—control invariance of the initial set and differentiability of its boundary—for ensuring that FRTs are themselves control invariant. These examples are based on a simple double integrator system defined by $\dot{p}=v,\dot{v}=u$, with state $x=[p\;v]^{T}$, and control input $u$, with control bound $u\in[-1,1]$. Note that curves $p=c\pm\frac{1}{2}v^{2}$ characterize trajectories that decelerate or accelerate until $v=0$, with the saturated input, $u=\pm 1$. The four specific $S$ sets are:
  1. (a)
    $S_{a}$ is defined as the circular region with radius $r$ centered at the origin. This set $S_{a}$ satisfies Assumption 2 but is not control invariant.
  2. (b)
    $S_{b}$ is formed by five curves, $p=-p_{1}+\frac{1}{2}v^{2}$, $p=p_{1}-\frac{1}{2}v^{2}$, $p=p_{2}-\frac{1}{2}v^{2}$, $p=-p_{2}+\frac{1}{2}v^{2}$, $p=-\frac{1}{2}v^{2}$, and the $p$-axis, as shown in Figure 2(b). This set does not satisfy Assumption 2, but is control invariant.
  3. (c)
    $S_{c}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$, and $p=p_{1}-\frac{1}{2}v^{2}$. This set also does not satisfy Assumption 2 but is control invariant.
  4. (d)
    $S_{d}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$ and $p=p_{1}-\frac{1}{2}v^{2}$ ($p_{1}>1$), and two arcs whose radius is $r=-1+2\sqrt{p_{1}}$ that are tangential to the curves whose centers are positioned at $(-p_{1}+r,0),(p_{1}-r,0)$, respectively. This set satisfies Assumption 2 and is also control invariant.
The first case demonstrates that the FRT can be a strict superset of $\mathrm{Int}(S_{a})$ when $S_{a}$ is not control invariant. The resulting $\mathrm{FRT}(\mathrm{Int}(S_{a}))$ is still not control invariant since the trajectory is bound to escape the set at points A and B. This example reveals a challenge in constructing a control invariant set with forward reachability when the initial set is not control invariant. In the second and third cases, the control invariance of $S$ can be checked analytically. The second case shows that the FRT can be a strict superset of $\mathrm{Int}(S_{b})$ if Assumption 2 is not met, even though $S_{b}$ is control invariant. Note that point C is where $f(x,\bar{\pi}(x))\!\in\!T_{S}(x)$ holds but $-f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(S)^{c}}(x)$ does not hold. In the third case, since $-f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(S)^{c}}(x)$ holds at both points D and E where $\partial S_{c}$ is not smooth, $\mathrm{FRT}(\mathrm{Int}(S_{c}))=\mathrm{Int}(S_{c})$. Finally, the set $S_{d}$ in the last case satisfies Assumption 2 and is also control invariant. Thus, according to Theorem 1, $\mathrm{FRT}(\mathrm{Int}(S_{d}))$ remains the same as $\mathrm{Int}(S_{d})$.
4 FRT value function and CBF
In this section, by taking the Hamilton-Jacobi approach to the forward reachability problem, we pose the computation of FRT as a differential game. The FRT is characterized by the value function proposed in Section 4.1. This value function is the unique solution to the HJ-PDE proposed in Section 4.2. In Section 4.3, under the assumption that the initial set is control invariant and has a differentiable boundary, we establish a connection between the proposed value function and the CBFs. Importantly, this provides an interpretation of any valid CBF as a forward reachability value function.
4.1 FRT Value Function with Discount Factor
By noting that $h_{S}$ satisfying Assumption 3 serves as a distance-like metric to the boundary of $S$ and its sign serves as an indicator of the inclusion in $S$, we can rewrite the definition of the FRT in (14) as follows:
$\mathrm{FRT}(\mathrm{Int}(S))$
=$=\!\big{\{}x~{}|~{}\exists\xi_{d}^{-}\in\Xi_{d}^{-},\forall\mathrm{u}^{-}\!\in\!\mathcal{U}^{-},\!\sup_{t\in(-\infty,0]}h_{S}(\mathrm{x}^{-}(t))\!>\!0\big{\}}$
=$=\!\big{\{}x~{}|~{}\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}h_{S}(\mathrm{x}^{-}(t))>0\big{\}}.$
Since rescaling $h_{S}(\mathrm{x}^{-}(t))$ with a positive constant at any time $t$ does not change its sign, the following holds:
$\mathrm{FRT}(\mathrm{Int}(S))\!=\Big{\{}x~{}|\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}^{-}(t))>0\Big{\}},$
where $\gamma\!>\!0$ and at each time $t\!\in\!(-\infty,0]$, $h_{S}(\mathrm{x}^{-}(t))$ is rescaled by $e^{\gamma t}$. Thus, by defining the FRT value function of $S$, $V_{\gamma}:\!\mathbb{R}^{n}\!\rightarrow\!\mathbb{R}$, as
$V_{\gamma}(x)\coloneqq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x,\mathrm{u}^{-},\Xi_{d}^{-})\vspace{-1em}$ (18)
$J_{\gamma}(x,\mathrm{u}^{-},\xi_{d}^{-})=\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}^{-}(t)),\vspace{-1em}$ (19)
where $\mathrm{x}^{-}$ solves (13) and $x$ is the terminal state of $\mathrm{x}^{-}$, the following holds.
Lemma 7.
Suppose $S\in\mathbb{R}^{n}$ is a closed set, $f$ satisfies Assumption 4, and a bounded function $h_{S}$ satisfies Assumption 3-1). $V_{\gamma}(x)$ is positive if and only if $x$ belongs to the FRT of the interior of $S$:
$\mathrm{FRT}(\mathrm{Int}(S))=\{x~{}|~{}V_{\gamma}(x)>0\}.$ (20)

### v4

@@ 2.2 Control Barrier Functions
Lemma 2 implies that a safe control input on the boundary of the set $S$ can render the trajectory $\mathrm{x}$ forward invariant in $S$. Control barrier functions (CBFs), first introduced in [4], additionally impose conditions on the input when the trajectory is strictly inside the set before reaching the boundary. Here, we extend its definition to robust CBF for the dynamics with disturbance (8).
@@
A function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Condition 2 for a closed set $S$ is a robust CBF for the dynamics (8) if there exists an extended class $\mathcal{K}$ function $\alpha$ such that, for all $x\in S$,
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)+\alpha\left(h_{S}(x)\right)\geq 0.\vskip-5.0pt$ (11)
@@
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)+\gamma h_{S}(x)\geq 0.\vskip-5.0pt$ (12)
This is the most common choice of class $\mathcal{K}$ function used in the CBF literature, and enables us to make a connection between CBFs and reachability value functions where $\gamma$ will play the role of a discount factor in the reachability formulation. Intuitively, (12) ensures that $h_{S}(\mathrm{x}(t))$ does not decay faster than the exponentially decaying curve $\dot{h}_{S}(\mathrm{x}(t))=-\gamma h_{S}(\mathrm{x}(t))$. This induces the braking mechanism to any trajectory $\mathrm{x}$ approaching the boundary of $S$. At $x\in\partial S$ where $h_{S}(x)=0$, (12) implies (10); thus, the existence of the robust CBF $h_{S}$ is a sufficient condition for $S$ being robust control invariant:
@@
If a robust CBF $h_{S}$ exists, $S$ is robustly control invariant.
@@
Since a robust CBF $h_{S}$ must satisfy Condition 2 (originally introduced in [4]), the invariant set $S$ must satisfy Condition 1 by Proposition 1.
2.3 Backward reachability for control invariance
We now introduce how backward reachability analysis is used to construct a maximal robust control invariant set contained in the constraint set $X\subset\mathbb{R}^{n}$. We consider the following definition of the backward reachable tube (BRT):
Definition 4 (Inevitable BRT).
For a given terminal set $C\subset\mathbb{R}^{n}$ which is an open set, we define the (infinite-horizon inevitable) BRT of $C$ as the following set.
$\mathrm{BRT}(C)\coloneqq\Big\{x\!\in\!\mathbb{R}^{n}~\Big|~\exists\xi_{d}\in\Xi_{d},T>0\text{ s.t.}\forall\mathrm{u}\in\mathcal{U},\;\;\;\;\;\;$
[…]
Consider the inevitable BRT of the failure region $X^{c}$. Then, the following holds:
Proposition 3.
For a closed set $X$, $\{\mathrm{BRT}(X^{c})\}^{c}$ is the maximal robust control invariant set contained in $X$ [12, 21].
This maximal robust control invariant set is also known as the (robust) viability kernel of the constraint set $X$ [5].
The $\mathrm{BRT}(X^{c})$ and its complement can be computed by considering the following BRT value function:
Definition 5 (BRT value function with discount factor [2, 49]).
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}e^{-\lambda t}h_{X}(\mathrm{x}(t)),$ (14)
where $h_{X}$ is a function that satisfies $X=\{x\in\mathbb{R}^{n}~|~h_{X}(x)\geq 0\}$. The control tries the maximize the minimum value of $h_{X}$ discounted in time. The value function is the viscosity solution [8] to the following Hamilton-Jacobi backward reachable tube variational inequality (HJ-BRT-VI):
$\displaystyle 0=\min\Big\{\$$h_{X}(x)-V(x),$ (15)
$\max_{u}\min_{d}\frac{\partial V}{\partial x}\cdot f(x,u,d)-\lambda V(x)\Big\}\vskip-7.5pt$
We now discuss why this value function has little relevance to robust CBFs in Definition 3, and is not ideal for safety control although it is useful to verify the maximal control invariant set. When $\lambda=\gamma>0$, because of the discount term applied to $h_{X}$ indefinitely in time to, for all states within the invariant set, $\inf_{t\in[0,\infty)}e^{-\lambda t}h_{X}(\mathrm{x}(t))$ flattens to zero, forming the maximal control invariant set as $\{x\in\mathbb{R}^{n}~|~V(x)=0\}$ [49] (Table 1, first row). Within the invariant set, the value function has zero gradient, making it not usable for constraining the control input as in the barrier constraint (12). In fact, as seen in (15), the value function satisfies the inequality condition
$\max_{u}\min_{d}\frac{\partial V}{\partial x}\cdot f(x,u,d)-\lambda V(x)\geq 0.\vskip-5.0pt$
To make this condition match with the barrier constraint, we need to consider $\lambda=-\gamma<0$, resulting in an anti-discount (or an exploding factor) to the cost function. The work in [13] presents this formulation for the finite-horizon BRT, which is named as the control barrier-value function (CBVF). However, its extension to infinite horizon might lead to an unbounded and discontinuous value function and non-unique solutions to the corresponding HJ-PDE (Table 1, second row). Similarly, the value function without any discount ($\lambda=0$) can lead to discontinuous and non-unique solutions to the corresponding HJ-PDE (Table 1, third row).
Finally, in general, the maximal control invariant set contained in the desired safety constraint set $X$ might have a non-differentiable boundary [5]. However, the differentiability of the boundary of the zero-superlevel set of the CBF is a feature derived from its definition, as discussed in Remark 1, which prohibits designing a CBF based on the maximal control invariant set.
These limitations of backward reachability approaches in designing CBFs motivate us to investigate the forward reachability as an alternative, which we discuss next.
@@ 3 Forward Reachability Analysis
We now introduce forward reachability analysis to control invariant sets. We first provide background on the forward reachability of a set.
@@ 3.1 Forward Reachability
Forward reachability analyzes a set’s evolution in the future. The forward reachable tube (FRT) of a set, roughly speaking, encompasses states that are reached by trajectories that depart from the initial set. This concept is illustrated in Figure 1. For autonomous systems (e.g. $\dot{\mathrm{x}}(t)=f(\mathrm{x}(t))$), the forward evolution of a set is uniquely determined (Fig 1 (a)). However, for systems with control and/or disturbance inputs, the trajectory, and thus the forward reachable tube, can be determined in various ways according to control and disturbance.
Consider the dynamics without disturbance, satisfying (1). At one extreme, the control can use its best effort to get further away from the original set $C$, and at the other extreme, the control works to stay as close to $C$ as possible. The former would cause the FRT to expand maximally covering all the states such that from $C$, reaching them is viable (Fig 1 (b)), and the latter would induce the FRT to grow minimally, encapsulating only the states which inevitably must have evolved from $C$ (Fig 1 (c)). From this intuition, we are able to define the viable (maximal) and inevitable (minimal) FRTs of a set $C$. Here, we only define the inevitable FRT which is the focus of this study, and readers are referred to [34] for the definition of the viable FRT. Applications of forward reachability in the safe control and verification literature primarily focused on determining the viable FRTs, and checking whether this set intersects with the unsafe set [3, 28, 48], with limited interests in the utility of inevitable FRTs [34].
@@
where $\mathrm{u}^{-}:(-\infty,0]\rightarrow U$ is an element of $\mathcal{U}^{-}$, a set of measurable backward control signal. Evaluating whether $\mathrm{x}^{-}$ reaches the set $C$ when time is considered to flow backward will tell us whether $x$ belongs to the forward reachable tube of $C$. With $\mathcal{D}^{-}$ denoting a set of measurable backward disturbance signals $\mathrm{d}^{-}:(-\infty,0]\rightarrow D$, $\xi_{d}^{-}$ is a non-anticipative strategy for the disturbance backward in time. The separate notation is necessary due to the causality of the disturbance strategy being reversed in time.
@@
Note that  $\mathrm{FRT}(\cdot)$ can be interpreted as a set mapping, $\text{FRT}:2^{\mathbb{R}^{n}}\rightarrow 2^{\mathbb{R}^{n}}$. In words, $\mathrm{FRT}(C)$ is a collection of states such that every trajectory reaching it forward in time must have passed through $C$ at some point in the past. The FRT is shaped by the control aiming to restrain the growth of the FRT, whereas the disturbance is assumed to act adversarially and attempts to grow the FRT. For control systems, the inevitable FRT is minimal [34] since it excludes any state that can be reached by a trajectory that does not evolve from $C$.
Figure 2: Forward reachable tubes under double integrator dynamics for various shapes of the initial set $C$. In the first row, $C$ is visualized as the interior of the pink level curve. The interior of the blue level curve in the second row is $\mathrm{FRT}(\mathrm{Int}(C))$ for each case. (a) Smooth $C_{a}$ that is not control invariant, resulting in $\mathrm{FRT}(\mathrm{Int}(C_{a}))\neq\mathrm{Int}(C_{a})$. The FRT is still not control invariant. (b, c) Nonsmooth sets $C_{b}$, $C_{c}$ that are control invariant. In case (b), $\mathrm{FRT}(\mathrm{Int}(C_{b}))\neq\mathrm{Int}(C_{b})$; this shows that Condition 1 is required for Corollary 1 to hold. In case (c), $\mathrm{FRT}(\mathrm{Int}(C_{c}))=\mathrm{Int}(C_{c})$. (d) Smooth control invariant $C_{d}$ that results in $\mathrm{FRT}(\mathrm{Int}(C_{d}))=\mathrm{Int}(C_{d})$ according to Corollary 1.
@@ 3.2 Forward Reachable Tubes of Control Invariant Sets
We now investigate the relationship between the inevitable FRTs and control invariance. Since the inevitable BRT of $X^{c}$ led to the discovery of the maximal control invariant set in the constraint set $X$ (Proposition 3), one might hypothesize that the inevitable FRT may lead to the minimal control invariant set containing the initial set $C$. However, we show that this is only true when the boundary of the FRT is differentiable:
@@
For any given set $C$, if $\mathrm{FRT}(\mathrm{Int}(C))$ satisfies Condition 1, $\mathrm{FRT}(\mathrm{Int}(C))$ is (the minimal) robust control invariant set under (8) containing $C$.